alexgame:

Solve for x: negative 6 over 5, multiplied by x minus 4 equals negative 46

SmokeyBrown:

\[-6/5 * ( x-4) = -46\] Is the equation we start with, based on this info. In the end, we want to get "x" by itself on one side of the "=". So, we can start by dividing both sides by (-6/5). This will give us: \[x-4 = 230/6\] Now we just have to add 4 to both sides, and we get \[x = (230/6) - 4\] And you can simplify that down if you want, but that'd pretty much be your answer

alexgame:

so its going to be 35

SmokeyBrown:

@alexgame wrote:
so its going to be 35
Pretty close! I got 34 1/3, which is right next to 35.